# Improving Human-AI Coordination through Adversarial Training and Generative Models

### v2

@@ 4 GOAT: Generative Online Adversarial Training
We introduce GOAT, which combines generative modeling with online adversarial training to optimize regret-based objectives and improve human-AI coordination.
Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods (Charakorn et al., 2023). Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it. This type of optimization is widely known as minimax optimization in game theory.
@@
As may be obvious, attempting to train agents to cooperate via XP minimization leads to a fundamentally misaligned optimization objective because it creates a conflict between agents to achieve the collective goal of maximizing common payoff. The Adversary agent is incentivized to develop behaviors that intentionally sabotage interactions rather than generate meaningful cooperation trajectories.
Generating adversarial cooperative agents. To address the above issues, we propose adding a pre-trained generative model (VAE) with frozen weights into the adversarial training loop. We hypothesize that adding a generative model that is trained on simulated agent coordination trajectories will ensure that adversarial training generates valid cooperative partners with no self-sabotage or handshaking issues. Previous XP minimization methods (Charakorn et al., 2023; Sarkar et al., 2023) represented policies as distinct, qualitative behaviors, requiring the algorithm to explore a combinatorial space of policies, whereas the VAE encodes the cooperation strategies as a continuous latent space allowing a smooth optimization landscape. We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation. Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce $z^{\prime}=\pi_{A}(z)$. This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$. Thus, the optimization problem is as follows:
@@
Note that this objective resolves the issue of directly training the partner policy that is choosing actions to minimize XP performance; Equation 3 reveals that the adversary agent does not have direct control over low-level actions, which instead are chosen by the frozen, pre-trained generative model. Instead, the adversary simply searches over the embedding space of the generative model to find challenging partners.
However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns. Because the generative model is capable of simulating low-scoring, poor-quality partners, since these exist in the training data, the minimax objective will encourage the Adversary to focus on sampling these agents to minimize XP score. Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies, which the Cooperator policy is not able to perform well with.
Regret-based adversarial cooperative training. To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent to learn diverse strategies. We now formalize the terms for cooperative settings that we will use to formulate Regret.
Self-Play (SP): This play style, first used by Samuel (1959) to play checkers, is defined as the expected return when an agent is paired with itself:
@@
Regret ($R$): As proposed initially by Robbins (1952), regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight. We formalize regret in the cooperative setting as the difference between the optimal self-play SP performance a policy can achieve with its training partner and the cross-play XP performance it achieves with another agent.
@@
  Training VAE on trajectory dataset $D={\tau_{i}}_{i=1}^{N}$ where $\tau={s_{i},a_{i},r_{i}}_{i=1}^{N}$
@@
At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$. This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$. The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns. Regret is evaluated after collecting the returns from both the game play styles. The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance. The full objective of the GOAT game is thus:
$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
The online, adversarial nature of this game ensures the Adversary is incentivized to continually challenge the Cooperator agent by discovering new partners it cannot play with yet, while the Cooperator is in turn corrects those deficiencies. As it does so, the regret decreases, and the Adversary is encouraged to explore the latent space to generate more challenging partners to maximize the regret. This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones. The GOAT is naturally incentivized to generate meaningful challenges. For invalid tasks, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no training signal to the Adversary. This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
@@ 5 Experiments
In this section, we will describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
@@
Figure 2: Carroll et al. (2020) first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit. To increase coordination complexity, Liang et al. (2024) introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup. This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
@@
Evaluation environment. We evaluate GOAT using the popular coordination benchmark Overcooked (Carroll et al., 2020), which requires two players to collaborate by dividing cooking tasks and avoiding obstructing each other. Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively. We test GOAT on the first five layouts introduced in Carroll et al. (2020), and add a more complex layout introduced by Liang et al. (2024), the Multi-Strategy Counter (all shown in Figure 2). Agents must follow small, intricate steps to prepare a dish successfully.
GOAT implementation. The Cooperator is trained using PPO, whereas the Adversary is method-independent, so we use REINFORCE to train it. Further details on the KL-regularization of the VAE and other architecture and hyperparameter details are available in the Appendix.
Baselines. We compare to prior baselines on human-AI coordination that are closely related to GOAT or are the previous state-of-the-art (SOTA). In particular, we evaluate GOAT against the conventional Behavior Cloning (BC) based method (Carroll et al., 2020) (BC+RL), population-based methods Fictitious Co-Play (FCP) (Strouse et al., 2022) and maximum entropy population (MEP) (Zhao et al., 2022), and GAMMA trained on MEP (MEP+GAMMA) Liang et al. (2024), which previously showed SOTA performance in real human evaluations.
@@
Human Evaluation. We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better. Because human studies are expensive, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter. The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans. Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol. Each participant was tasked with completing 6 total rounds of the Overcooked game with an AI agent partner, where each round presented a different agent model among one of the baselines or our method. In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds. Additionally, each participant is instructed to answer qualitative questions on a 5-point Likert scale in order to rate their experience playing with each agent partner.
@@ 6 Experimental Results
Below, we described the results of empirical experiments organized around our central research questions.
@@
RQ1: Will adversarial training result in a more robust Cooperator with better sample complexity? PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors. Training a Cooperator against them may not achieve good performance with people, and it may take a long time since the training does not focus on correcting weaknesses in the Cooperator but instead randomly samples from the population. We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time.
@@
To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 3. Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work (Carroll et al., 2020; Strouse et al., 2022; Liang et al., 2024). The human proxy enables the automatic evaluation of agents’ performance throughout training.
Figure 3: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC). We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method (Liang et al., 2024). Error bars show the standard error over 5 random seeds. The Adversary is able to achieve better performance in fewer samples (RQ1).
@@
As is evident in the curves, our adversarial training method achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques. We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations. Adversarial exploration of this space should enable rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies. By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to real human partners. Because it focuses training on weak points in the Cooperator’s policy, it creates an agent that is more robust to human stratgies, with lower sample complexity than traditional PBT methods.
CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
MEP $1.0$ $3.48\times 10^{-28}$
MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
[…]
@@
RQ2: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans? As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts. We recruited novel human players whom the agents had not previously seen. Figure 4 depicts the evaluation results for the layout and all the agents against human players. Each entry corresponds to a human-agent team. On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA see Table 1. We hypothesize this is because both methods are reaching close to optimal performance on this layout. However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art.
@@
Figure 4: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit and multi-strategy counter.
RQ3: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective? In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate. In Equation 3, we showed how integrating a generative model could help alleviate this problem. However, while the generative model can prevent sabotage behavior, that does not mean all agents it simulates will show productive cooperation behavior. We claimed that the purely adversarial minimax objective would still focus too much on poorly performing or incompetent partners, motivating the use of the regret objective in Equation 5. We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout. As we can see in Figure 6, the GOAT significantly outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT.
We further investigate the embeddings chosen by both types of Adversary in Figure 6 throughout the training period. We find that the minimax Adversary is pushed to a region (Red cluster in the plot) where the policies result in low reward behaviors. One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance but makes no attempts to cooperate. In the blue, we have embeddings chosen by the GOAT throughout the training period. We can observe that the regret objective enables the GOAT to explore the latent space effectively, thus covering a wider range of interesting behaviors.
Figure 5: Evaluation curves for comparison of performance of minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner. Error bars are the Standard Error of the Mean.
Figure 6: Adversary in red and GOAT in blue exploring the latent space over the training period.
@@
RQ4: Will adversarial training create a curriculum of increasingly challenging strategies?
Figure 7: GOAT exploring the latent space exploration compared against GAMMA (Liang et al., 2024) in red over the whole training period.
@@
The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges. As GOAT identifies underperforming coordination strategies, it should generate increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors. This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods. Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
@@
The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 7. We compare this to the distribution of embeddings the original GAMMA method (Liang et al., 2024) samples throughout training (red ball). We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple). As training progresses, it chooses a particular region to exploit. After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret. Finally, at the end of the training step in yellow, we can see that the GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.

### v3

@@ 4 GOAT: Generative Online Adversarial Training
We introduce GOAT, which combines generative modeling with online regret-based adversarial training to enable robust human-AI coordination.
Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods [4]. Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it. This type of optimization is widely known as minimax optimization in game theory:
@@
As may be obvious, attempting to train agents to cooperate via XP minimization leads to a fundamentally misaligned optimization objective because it conflicts with the cooperative objective of maximizing common payoff. The Adversary agent is incentivized to develop behaviors that intentionally sabotage interactions rather than generate meaningful cooperation trajectories.
Generating adversarial cooperative agents. To address the above issues, we propose adding a pre-trained generative model (VAE) with frozen weights into the adversarial training loop. We hypothesize that adding a generative model trained solely on cooperative policy data will ensure that adversarial training generates only valid cooperative partners. By design, the generative model is not capable of simulating self-sabotage or handshaking behaviors. Instead, we train a separate Adversary policy to search for embedding vectors $z$ that decode into partners that minimize XP performance for the specific Cooperator agent we seek to train. The Adversary can thus search over a well-structured, continuous latent space, allowing a smooth optimization landscape.
We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation 1. Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce an Adversary embedding $z^{\prime}=\pi_{A}(z)$. This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$. Thus, the optimization problem is as follows:
@@
However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns. Therefore, the Adversary will only search for the worst agents that the generative model can simulate. Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies which could have good performance, but the Cooperator failed to cooperate well with them.
Regret-based adversarial cooperative training. To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent. We now formalize the definition of Regret we will use for cooperative settings.
Self-Play (SP): This play style, first used by Samuel [27] to play checkers, is defined as the expected return when an agent is paired with itself:
@@
Regret ($R$): As proposed initially by Robbins [25], regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight. We formalize regret in the cooperative setting as the performance gap between the cross-play (XP) performance an agent achieves with a partner $\pi_{P}$, and $\pi_{P}$’s optimal self-play (SP) performance.
@@
At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$. This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$. The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns. Regret is evaluated after collecting the returns from both games. The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance. The full objective of the GOAT game is thus:
$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
The online optimization as shown in algorithm 1 ensures that the Adversary is incentivized to continually discover new partner strategies that are challenging for the Cooperator, by searching the latent space to maximize regret. The Cooperator in turn learns to adapt to these partner strategies by minimizing the regret. This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones. For invalid partners, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no incentive to the Adversary. This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
@@
  Training VAE on trajectory dataset $D={\tau_{i}}_{i=1}^{N}$ where $\tau=\{s_{i},a_{i},r_{i}\}_{i=1}^{N}$
@@ 5 Experiments
In this section, we describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans. We evaluate GOAT using two popular coordination benchmarks: 1)1)1 ) One-Step Cooperative Matrix Game (CMG) [20, 43, 4] and 2)2)2 ) Overcooked [3].
@@
Figure 2: Carroll et al. [3] first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit. To increase coordination complexity, Liang et al. [18] introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup. This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
5.1 One-Step Cooperative Matrix Game (CMG):
This is a single-step game where two agents (Row agent and Column agent) independently select an action by choosing a row and a column, respectively, and the intersection of their choices returns the associated reward according to the payoff matrix. Both agents receive the same reward and attempt to maximize the common payoff. A typical payoff matrix in CMG is defined by $(M,\{k_{m}\},\{r_{m}\})$ where $M$ is number of solutions, $k_{m}$ is number of compatible actions, and $r_{m}$ is the reward associated with each solution. We use a challenging multimodal layout called CMG-M as shown in Figure 3.
5.2 Overcooked:
@@
This is a collaborative game where two players need to collaborate by dividing cooking tasks and avoiding obstructing each other. Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively. We test GOAT on the first five layouts introduced in [3], and add a more complex layout introduced by Liang et al. [18], the Multi-Strategy Counter (all shown in Figure 2).
5.3 GOAT implementation & Baselines:
The Cooperator is trained using PPO, and the Adversary is trained using REINFORCE. Training the Adversary is RL-algorithm-agnostic as it does not depend on state and is a one-step optimization process. Further details on the KL-regularization of the VAE, training procedure for GOAT, and other architecture and hyperparameters are available in the Appendix. We compare to 5 competitive prior baselines on human-AI coordination: Behavior Cloning (BC) partners with RL cooperator [3] (BC+RL), PBT methods Fictitious Co-Play (FCP) [32], maximum entropy population (MEP) [43], CoMeDi [28], and GAMMA trained on MEP (MEP+GAMMA) [18], which previously showed state-of-the-art performance in real human evaluations.
5.4 Human Evaluation:
@@
We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better. Because we want the most challenging tasks with the most diverse strategy space, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter. The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans. Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol. Each participant was tasked with completing 6 total rounds of the Overcooked game with an AI agent partner, where each round presented a different agent model among one of the baselines or our method. In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds.
@@ 6 Experimental Results
(a) CMG Layout
(b) Self-play
(c) TrajeDi
(d) MEP
(e) CoMeDi
(f) GOAT
Figure 3: To test the ability of the method to discover and exploit multiple minima and maxima in the problem space, we introduce a custom CMG layout (a). This reward structure is designed to have three scenarios: no reward, suboptimal rewards, and global maximum reward. The policies for the methods were aggregated and normalized into a heatmap.
6.1 RQ1: Can GOAT discover and exploit multiple cooperative modes?
For GOAT to be effective, the VAE should be able to discover and exploit regions of the reward space. To test this, we use a one-step cooperative matrix game (CMG). We test the cooperative policy generation capability of VAE by training it on trajectory data generated by the population of CoMeDi agents. Since CoMeDi data is cooperative, we hypothesize that the VAE should only generate policies that are in distribution. To keep the evaluation fair, we used a population size of 8 for all the methods tested on CMG-M. During training, GOAT covers all the modes and correctly assigns importance to the agent’s actions. As seen in Figure 3, the baselines self-play, Trajedi, and MEP get stuck and might not discover other modes. CoMeDi discovers all the regions and gives the highest priority to the global maximum, but spreads probability mass roughly equally across all other modes, regardless of their payoff. In contrast, GOAT not only covers all modes but is better at assigning higher probability to more rewarding modes (for example, assigning low probability to modes 4 and 7, which have low payoff).
We note that the quality of VAE training and the amount of training data greatly influence the resulting GOAT policy. To robustly train the VAE, we follow Liang et al. [18] and augment the 8 policy pool data by pairing the row and column agents of one policy with other policies. In our experiments, we saw that higher $\beta$ for KL regularization in VAE resulted in better generation of actions.
6.2 RQ2: Will adversarial training result in a more robust Cooperator with better sample complexity?
@@
PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors. Training a Cooperator against them may not achieve good performance with people, and it may take a long time since the training does not focus on correcting weaknesses in the Cooperator but instead randomly samples from the population. We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time.
@@
To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 4. Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work [3, 32, 18]. The human proxy enables the automatic evaluation of agents’ performance throughout training.
@@
As is evident in the curves, GOAT achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques. We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations. Adversarial exploration of this space enables rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies. By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to diverse partners. Because it efficiently targets training to weak points in the Cooperator’s policy, it has lower sample complexity than traditional PBT methods.
Figure 4: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC). We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method, GAMMA [18]. Error bars show the standard error over 5 random seeds.
6.3 RQ3: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans?
@@
As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts. We recruited novel human players whom the agents had not previously seen. Figure 5 depicts the evaluation results for the two layouts and all the agents against human players. On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA (see Table 4 for statistical significance results). We hypothesize this is because both methods are reaching close to optimal performance on this layout. However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art. To test these differences for yourself, please visit our webpage (https://sites.google.com/view/goat-2025/), which provides an interactive demo in which you can play Overcooked with both GOAT and the baselines.
@@
Figure 5: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit (left plot) and multi-strategy counter (right plot).
@@
6.4 RQ4: Will adversarial training create a curriculum of increasingly challenging strategies by discovering them in latent space?
@@
The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges. GOAT identifies underperforming coordination strategies and generates increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors. This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods. Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
Figure 6: Left: GOAT exploring the latent space exploration compared against GAMMA [18] in red over the whole training period. Right: Minimax Adversary in red and GOAT in blue exploring the latent space over the training period. We provide a visual example related to this on our webpage.
@@
The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 6. We compare this to the distribution of embeddings of the original GAMMA method [18] samples throughout training (red ball). We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple). As training progresses, it chooses a particular region to exploit. After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret. Finally, at the end of the training step in yellow, GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
6.5 RQ5: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective?
In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate. In Equation 3, we showed how integrating a generative model could help alleviate this problem. However, while the generative model can prevent sabotage behavior, that does not mean all agents it simulates will show productive cooperation behavior. We claimed that the purely adversarial minimax objective would still focus too much on poorly performing or incompetent partners, motivating the use of the regret objective in Equation 5. We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout. As we can see in Figure 7, the GOAT outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT. We further investigate the embeddings chosen by both types of Adversary throughout the training period in Figure 6 (right). We find that the minimax Adversary mostly fixates on a single region (red cluster in the plot) where the policies result in low reward behaviors. One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance, but makes no attempts to cooperate. In the blue, we have embeddings chosen by the GOAT throughout the training period. We can observe that the regret objective causes the GOAT to move to explore multiple modes over the course of training, exploring the latent space more effectively and thus covering a wider range of interesting behaviors.